OpenStudy (missykea):

Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (anonymous):

1,3

OpenStudy (anonymous):

Have you done anything for this?

OpenStudy (texaschic101):

-2x + y = 1 -4x + y = -1 --- (-1) -------------- -2x + y = 1 4x - y = 1 (result of multiplying by -1) --------------add 2x = 2 -- divide by 2 on both sides x = 1 now sub 1 in for x in either equation to find y -2x + y = 1 -2(1) + y = 1 -2 + y = 1 -- add 2 to both sides y = 1 + 2 y = 3 check... -4x + y = -1 -4(1) + 3 = -1 -4 + 3 = -1 -1 = -1 (correct) solution is (1,3)